OpenStudy (anonymous):

Solve for the indicated variable. P = 3r + 2s; for s S=

OpenStudy (anonymous):

well S and s are different lol

OpenStudy (anonymous):

so S := obama XD

OpenStudy (the_fizicx99):

^^ P = 3r + 2s -3r -3r p - 3r = 2s /2 /2 P - 3r/2 = S OR S = P -3r/2